OpenStudy (anonymous):

The books in a private library are classified as fiction and nonfiction. there are 400 books in the library. there are 40 more fiction that nonfiction books. audrey randomly picks a book a few minutes later . ryan randomly picks one of the remaining books. what is the probability that both pick nonfiction books?

OpenStudy (anonymous):

First let's find how many of each type of book there are. f = Fiction books nf = Non-Fiction book f = nf+40 f+nf=400 - Substitute "f" for (nf+40) - nf+40+nf = 400 To solve for "nf" we have to get rid of any other distractions, to do that we subtract 40 from both sides 2nf = 360 nf = 180 f = 220 Now we know how many of each type of books there are 180 Non-Fiction books 220 Fiction books First pick (Total Non-Fiction books/Total books) = 180/400 Second pick (Total Non-Fiction books remaining/Total books remaining) = 179/399 Since the question asks for Ryan *and* Audrey, we take the 2 probabilities and times them together. This gets us 32220/159600

OpenStudy (anonymous):

You can simplify the fraction further...